# SelectIT: Selective Instruction Tuning for Large Language Models via Uncertainty-Aware Self-Reflection

### v1

@@ Abstract
Instruction tuning (IT) is crucial to tailoring large language models (LLMs) towards human-centric interactions. Recent advancements have shown that the careful selection of a small, high-quality subset of IT data can significantly enhance the performance of LLMs. Despite this, common approaches often rely on additional models or data sets, which increases costs and limits widespread adoption. In this work, we propose a novel approach, termed SelectIT, that capitalizes on the foundational capabilities of the LLM itself. Specifically, we exploit the intrinsic uncertainty present in LLMs to more effectively select high-quality IT data, without the need for extra resources. Furthermore, we introduce a novel IT dataset, the Selective Alpaca, created by applying SelectIT to the Alpaca-GPT4 dataset. Empirical results demonstrate that IT using Selective Alpaca leads to substantial model ability enhancement. The robustness of SelectIT has also been corroborated in various foundation models and domain-specific tasks. Our findings suggest that longer and more computationally intensive IT data may serve as superior sources of IT, offering valuable insights for future research in this area. Data, code, and scripts are freely available at https://github.com/Blue-Raincoat/SelectIT.
@@ 1 Introduction
Large language models (LLMs) have attracted much attention due to their impressive capabilities in following instructions and solving intricate problems (Touvron et al., 2023b, a; Achiam et al., 2023; Penedo et al., 2023). A crucial aspect of enhancing LLMs’ performance is instruction tuning (IT), which involves the supervised adjustment of LLMs using pairs of instructional data, essential for refining the models’ ability to accurately respond to human instructions. Recent groundbreaking research, such as the LIMA Zhou et al. (2023), highlights the critical importance of instructional data quality over quantity. Contrary to the approach of merely increasing the dataset size, a carefully selected, smaller dataset of higher quality can significantly improve LLMs’ performance.
Figure 1: Existing advanced data selection strategies rely heavily on external models or data sets; however, SelectIT effectively overcomes this limitation.
Despite the emergence of various methods for selecting high-quality data, these methods often rely on external resources and limit broader implementation. External ModelChen et al. (2023a); Liu et al. (2023) propose the employment of closed-source LLMs to evaluate or rank IT data. To circumvent the closed-source limitations, Li et al. (2023a, b); Kung et al. (2023) recommend the adoption of fine-tuned open-source LLMs, which will necessitate increased computational resource expenditure. External DataCao et al. (2023) split all mixed data into several bins and fully trained the models to evaluate different indicators of high-quality IT data. Despite these advancements, the challenge of conducting precise and efficient high-quality data selection without external resources remains unresolved.
We introduce SelectIT, a novel approach designed to enhance IT data selection by fully leveraging the foundation model itself, eliminating the need for external resources. SelectIT employs different grain uncertainty of LLMs: token, sentence, and model, which can effectually improve the accuracy of IT data selection. We first use the foundation model itself to rate the IT data from 1 to $K$ based on the uncertainty of various tokens. Next, we use sentence-level uncertainty to improve the rating process by exploiting the effect of different prompts on LLMs. At a higher model level, we utilize the uncertainty between different LLMs, enabling a collaborative decision-making process for IT data selection. By applying SelectIT to the original Alpaca, we curate a compact and superior IT dataset, termed Selective Alpaca.
Experimental results show that SelectIT outperforms existing high-quality data selection methods, improving LLaMA-2’s performance on the open-instruct benchmark Wang et al. (2023). Further analysis reveals that SelectIT can effectively discard abnormal data and tends to select longer and more computationally intensive IT data. The primary contributions of SelectIT are as follows:
@@
  • We introduce a novel IT dataset, Selective Alpaca, by selecting the high-quality IT data from the Alpaca-GPT4 dataset.
@@ 2 Related Work
The recognition of IT data quality’s superiority over quantity in the context of instruction tuning is well-established, yet the efficient and precise identification of high-quality data continues to be a challenging frontier for research. One straightforward approach is utilizing the closed-source advanced LLMs for IT data evaluation and selection Chen et al. (2023a); Liu et al. (2023). To circumvent the constraints associated with closed-source, some researchers opt to fine-tune LLMs directly to select high-quality IT data Li et al. (2023b); Kung et al. (2023). Instruction Mining Cao et al. (2023), involves training multiple models to derive various metrics, using statistical regression to develop a data quality assessment framework. Li et al. (2023a) propose training a specialized model and utilizing two unique, condition-based losses on this for a comprehensive IT data selection. Additionally, Wu et al. (2023) explore where data selection is informed by the similarity of samples within the embedding space of a fine-tuned model. However, the methods described above depend, to varying degrees, on supplementary datasets, the use of closed-source models, or open-source models that have been specially fine-tuned, which results in increased consumption of resources and limits the broader impact.
Figure 2: The overall framework of SelectIT.
@@ 3 Our SelectIT Method
Utilizing advanced LLMs for the sample evaluation is a widely adopted approach in the IT data selection Chen et al. (2023a); Li et al. (2023b); Liu et al. (2023). Given an IT dataset $D$ containing a sample $S$ = (input $X$, response $Y$), a designated rating prompt $RP$, and the LLMs $M$, the goal is to leverage both $RP$ and $S$ to prompt $M$ to assign an evaluation $Score$ to the sample $S$ on a scale from 1 to $K$. A higher score typically signifies superior IT data $Quality$.
@@
While existing methods Chen et al. (2023a); Cao et al. (2023) are adept at identifying high-quality samples, they often over-rely on external resources. To address these challenges, we introduce SelectIT, a strategy that capitalizes on the internal uncertainty of LLMs to efficiently select high-quality IT data. SelectIT incorporates three grains of sample evaluation modules: token, sentence, and model-level self-reflections, which effectively improve the reliability of IT data selection. The comprehensive framework of SelectIT is depicted in Figure 2.
@@ 3.1 Token-level Self-Reflection
Numerous studies have demonstrated that foundation models exhibit robust capabilities for next-token prediction during their pre-training phase Touvron et al. (2023b, a). Yet, this predictive strength is frequently underutilized in evaluating IT data quality. In SelectIT, we adopt a similar idea to evaluate IT data. Specifically, we calculate the next-token (from 1 to $K$) probability based on the rating prompt $RP$ and sample $S$. The score token with the highest probability is then considered as the sample’s quality.
$S^{base}=\underset{k\in\{1,\ldots,K\}}{\arg\max}\,P^{\prime}_{k},P^{\prime}_{k}=\left(\frac{e^{P_{k}}}{\sum_{j=1}^{K}e^{P_{j}}}\right)$ (2)
where $P_{k}$ and $P^{\prime}_{k}$ mean the probability and softmaxed probability of token $k$. $K$ means the number of scores to be considered.
@@ 3.2 Sentence-level Self-Reflection
Different prompts can significantly affect outputs of LLMs Kung et al. (2023); Peng et al. (2023b), introducing uncertainty into IT data evaluation at the sentence level. To make better use of this uncertainty to bolster the reliability of our method, we implement sentence-level self-reflection (Sentence-R). Building upon Token-R, we devise $K$ semantically similar rating prompts$\{RP_{0},RP_{1},\ldots,RP_{K}\}$ to obtain a series of quality scores $\{S^{token}_{0},S^{token}_{1},\ldots,S^{token}_{K}\}$ based on a given sample $S$. We calculate the average of these scores to represent the overall quality of sample $S$, because of the importance of incorporating assessments from diverse prompts. Additionally, we use the standard deviation to quantify the LLMs’ uncertainty to rating prompt; a higher standard deviation suggests greater sensitivity to prompt variation, while a lower standard deviation indicates more consistent and confident quality ratings by LLMs (Zhou et al., 2020). By integrating a holistic sample evaluation with the quantification of model uncertainty, we derive the sentence-level score $S^{sent}$, offering a more nuanced and reliable measure of IT data quality.
$S^{sent}=\frac{\textbf{Avg}\{S^{token}_{i}\}^{K}_{i=1}}{1+\alpha\times{\textbf{Std}\{S^{token}_{i}\}^{K}_{i=1}}}$ (4)
@@ 3.3 Model-level Self-Reflection
A sample affirmed by multiple foundation models can truly be deemed as high-quality. Different foundation models have different quality assessments of the sample, which introduce model-level uncertainty. To maximize the utilization of this uncertainty, we introduce model-level self-reflection (Model-R). This strategy is in harmony with our principles, as it leverages the capabilities of existing open-source models without the need for additional resources or the complexities associated with fine-tuning However, the challenge lies in the diverse capabilities of various LLMs and determining how to reasonably combine their sample evaluation based on their performance. It is widely acknowledged that the capabilities of LLMs tend to increase with their parameter count Hendrycks et al. (2021). Thus, we suggest using the parameter count of LLMs as an initial metric for assessing their capabilities to properly weight sample quality scores. Given $N$ foundation models with parameter counts $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ and their respective sentence-level scores for a sample $S$ being $\{S^{sent}_{0},S^{sent}_{1},\ldots,S^{sent}_{N}\}$, we formulate the model-level score $S^{model}$ to reflect a comprehensive evaluation of sample quality.
@@
where $N$ means the number of the foundation models. We use $S^{model}$ as the final evaluation of sample $S$ in SelectIT. The higher $S^{model}$, the better sample quality. We sort the samples in descending order based on their $S^{model}$ and then select the top-ranked samples as high-quality data.
@@ 3.4 Selective Alpaca
We apply SelectIT to the widely-used Alpaca-GPT4 (Peng et al., 2023a). Specifically, we use the most popular LLaMA-2 (7B, 13B, 70B) as our foundation models and set the hyper-parameters $K=5$, which decides the range of LLMs rating in Token-R and the number of rating prompts in Sentence-R. We finally select the top 20% as the high-quality data and obtain a novel IT dataset called Selective Alpaca.
@@ 4 Experiments
ID   System External MMLU BBH GSM TydiQA CodeX AE Overall
Model Data AVG $\Delta$ (normal-$\uparrow$)
Base Model: LLaMA-2-7B Implemented Existing Method
 1    Alpaca-GPT4 46.2 39.2 15.0 43.3 27.8 33.7 34.3 -
 2    LIMA 45.8 38.5 14.0 45.5 24.2 33.1 33.5 -0.7
 3    1 + AlpaGasus 46.8 39.2 14.5 48.4 26.5 34.6 35.0 +0.8
 4    1 + Q2Q 46.7 39.8 16.5 45.5 28.1 35.1 35.3 +1.1
 5    1 + Instruction Mining 47.0 40.0 16.5 47.8 29.6 34.4 35.9 +1.7
[…]
 6    1 + Token-R 47.7 38.7 15.0 48.6 28.0 35.3 35.5 +1.3
 7    1 + Sentence-R 46.7 39.9 16.5 48.9 27.8 34.5 35.7 +1.5
 8    1 + Model-R 47.5 38.0 16.5 45.9 28.7 35.6 35.4 +1.2
[…]
 9    SelectIT (6 + 7 + 8) 47.7 40.8 17.5 47.8 29.8 35.4 36.5 +2.3
Base Model: LLaMA-2-13B Implemented Existing Method
10    Alpaca-GPT4 55.7 47.3 31.0 49.1 41.8 46.5 45.2 -
11    LIMA 54.6 48.3 29.5 51.2 34.4 43.6 43.6 -1.6
12   10 + AlpaGasus 54.1 49.3 32.0 52.6 39.3 47.5 45.8 +0.6
13   10 + Q2Q 55.3 48.5 34.0 50.8 42.3 47.7 46.4 +1.2
14   10 + Instruction Mining 55.5 49.7 33.0 51.2 41.5 46.1 46.2 +1.0
[…]
15   10 + Token-R 55.3 47.9 31.0 53.8 40.8 46.2 45.8 +0.6
16   10 + Sentence-R 55.2 48.5 34.0 54.2 41.5 47.3 46.8 +1.6
17   10 + Model-R 55.5 47.2 33.5 53.0 42.2 47.1 46.4 +1.2
[…]
18   SelectIT (15 + 16 + 17) 55.5 50.1 34.5 54.4 43.2 48.2 47.7 +2.5
Table 1: The overall results on IT. “Codex” and “AE” mean HumanEval and AlpacaEval benchmark respectively.
@@ 4.1 Setups
To gain a more comprehensive understanding of the capabilities of LLMs, we evaluate our approach in diverse downstream tasks Wang et al. (2023); Ivison et al. (2023). Factual knowledge: We use the Massive Multitask Language Understanding dataset (MMLU Hendrycks et al. (2021)) to assess the factual knowledge of LLMs and report 5-shot results. Reasoning: We evaluate the reasoning abilities of LLMs using two widely utilized datasets: the Grade School Math dataset (GSM (Cobbe et al., 2021)) and Big-Bench-Hard (BBH Suzgun et al. (2022)) with the CoT setting Wei et al. (2022b). Multilinguality: we assess this ability by TyDiQA, a multilingual question-answering benchmark that encompasses 11 diverse languages, with the gold-passage setup. Coding: We evaluate this ability using the HumanEval dataset (Chen et al., 2021) and report pass@10 results with a temperature of 0.8. Open-ended generation: We utilize AlpacaEval (Dubois et al., 2023), which employs GPT-4 to effectively assess model outputs. This can evaluate whether the text produced by LLMs aligns with humans.
@@
The compared baselines are:
@@
[…]
  • AlpaGasus (Chen et al., 2023a) involves utilizing the robust ChatGPT to score and select data from the original Alpaca dataset.

### v2

# SelectIT: Selective Instruction Tuning for LLMs via Uncertainty-Aware Self-Reflection

@@ Abstract
Instruction tuning (IT) is crucial to tailoring large language models (LLMs) towards human-centric interactions. Recent advancements have shown that the careful selection of a small, high-quality subset of IT data can significantly enhance the performance of LLMs. Despite this, common approaches often rely on additional models or data, which increases costs and limits widespread adoption. In this work, we propose a novel approach, termed SelectIT, that capitalizes on the foundational capabilities of the LLM itself. Specifically, we exploit the intrinsic uncertainty present in LLMs to more effectively select high-quality IT data, without the need for extra resources. Furthermore, we introduce a curated IT dataset, the Selective Alpaca, created by applying SelectIT to the Alpaca-GPT4 dataset. Empirical results demonstrate that IT using Selective Alpaca leads to substantial model ability enhancement. The robustness of SelectIT has also been corroborated in various foundation models and domain-specific tasks. Our findings suggest that longer and more computationally intensive IT data may serve as superior sources of IT, offering valuable insights for future research in this area. Data, code, and scripts are freely available at https://github.com/Blue-Raincoat/SelectIT.
@@ 1 Introduction
Figure 1: Existing advanced data selection strategies rely heavily on external models or data; however, SelectIT effectively overcomes this limitation.
Large language models (LLMs) have attracted much attention due to their impressive capabilities in following instructions and solving intricate problems (Touvron et al., 2023b, a; Achiam et al., 2023; Penedo et al., 2023). A crucial aspect of enhancing LLMs’ performance is instruction tuning (IT), which involves the supervised adjustment of LLMs using pairs of instructional data, essential for refining the models’ ability to accurately respond to human instructions. Recent groundbreaking research, such as the LIMA (Zhou et al., 2023), highlights the critical importance of instructional data quality over quantity. Contrary to the approach of merely increasing the dataset size, a carefully selected, smaller dataset of higher quality can significantly improve LLMs’ performance.
Despite the development of various high-quality data selection methods, they often depend on external resources, limiting wider implementation. External ModelChen et al. (2024); Liu et al. (2023) propose the employment of closed-source LLMs to evaluate or rank IT data. To circumvent the closed-source limitations, Li et al. (2023a, b); Kung et al. (2023) recommend fine-tuning open-source LLMs, which requires more computational resources. External DataCao et al. (2023) split all mixed data into several bins and fully trained the models to evaluate different indicators of high-quality IT data. Despite these advancements, the challenge of precise and efficient high-quality data selection without external resources remains unresolved.
In this paper, we introduce SelectIT, a novel approach designed to enhance IT data selection by fully leveraging the foundation model itself, eliminating the need for external resources. SelectIT employs different grain uncertainty of LLMs: token, sentence, and model, which can effectually improve the accuracy of IT data selection. We first use the foundation model itself to rate the IT data from 1 to $K$ based on the uncertainty of various tokens. Next, we use sentence-level uncertainty to improve the rating process by exploiting the effect of different prompts on LLMs. At a higher model level, we utilize the uncertainty between different LLMs, enabling a collaborative decision-making process for IT data selection. By applying SelectIT to the original Alpaca, we curate a compact and superior IT dataset, termed Selective Alpaca.
Experimental results show that SelectIT outperforms existing high-quality data selection methods, improving LLM’s performance on the open-instruct benchmark (Wang et al., 2024). Further analysis reveals that SelectIT can effectively discard abnormal data and tends to select longer and more computationally intensive IT data. The primary contributions of SelectIT are as follows:
@@
  • We introduce a curated IT dataset, Selective Alpaca, by selecting the high-quality IT data from the Alpaca-GPT4 dataset.
@@ 2 Related Work
The recognition of IT data quality’s superiority over quantity in the context of IT is well-established, yet the efficient and precise identification of high-quality data continues to be a challenging frontier for research. One straightforward approach is utilizing the closed-source advanced LLMs for IT data evaluation and selection (Chen et al., 2024; Liu et al., 2023). To circumvent the constraints associated with closed-source, existing research opt to fine-tune LLMs directly to select high-quality IT data (Li et al., 2023b; Kung et al., 2023). Li et al. (2023c); Gururangan et al. (2020); Chen et al. (2023a); Cao et al. (2023) use pre-defined notions of useful data or other IT datasets to develop a data quality assessment framework. Li et al. (2023a) propose training a specialized model and utilizing two unique, condition-based losses on this for a comprehensive IT data selection. Wu et al. (2023) explore where data selection is informed by the similarity of samples within the embedding space of a fine-tuned model. N-gram features (Xie et al., 2023) or model gradients (Xia et al., 2024; Han et al., 2023) are also important features for selecting high-quality data in fine-tuned LLMs. However, the methods described above depend, to varying degrees, on supplementary datasets, the use of closed-source models, or open-source models that have been specially fine-tuned, which results in increased consumption of resources and potentially limits the broader impact.
Figure 2: Overall framework of SelectIT. In Token-level Self-Reflection, we employ the foundation model to rate the IT data from 1 to $K$. In Sentence-level Self-Reflection, we leverage the uncertainty of varied prompts on LLMs to enhance the rating process. In Model-level Self-Reflection, we harness uncertainty among different LLMs to facilitate a collaborative decision-making process in selecting IT data. Finally, different levels of self-reflection are reasonably combined into SelectIT, which can effectively select high-quality IT data without relying on additional resources.
@@ 3 Our SelectIT Method
Utilizing advanced LLMs for the sample evaluation is a widely adopted approach in the IT data selection (Chen et al., 2024; Li et al., 2023b; Liu et al., 2023). Given an IT dataset $D$ containing a sample $S$ = (input $X$, response $Y$), a designated rating prompt $RP$, and the foundation LLMs $M$, the goal is to leverage both $RP$ and $S$ to prompt $M$ to assign an evaluation $Score$ to the sample $S$ on a scale from 1 to $K$. A higher score typically signifies superior IT data $Quality$.
@@
While existing methods (Chen et al., 2024; Cao et al., 2023) are adept at identifying high-quality samples, they often over-rely on external resources. To address these challenges, we introduce SelectIT, a strategy that capitalizes on the internal uncertainty of LLMs to efficiently select high-quality IT data. SelectIT incorporates three grains of sample evaluation modules: token, sentence, and model-level self-reflections, which effectively improve the reliability of IT data selection. The comprehensive framework of SelectIT is depicted in Figure 2.
@@ 3.1 Token-level Self-Reflection
Numerous studies have demonstrated that foundation models exhibit robust capabilities for next-token prediction during their pre-training phase (Touvron et al., 2023b, a). Yet, this predictive strength is frequently underutilized in evaluating IT data quality. In SelectIT, we adopt a similar idea to evaluate IT data. Specifically, we calculate the next-token probability (from 1 to $K$) based on the rating prompt $RP$ and sample $S$. The score token with the highest probability is then considered as the sample’s quality.
$S^{base}=\underset{k\in\{1,\ldots,K\}}{\arg\max}\,P^{\prime}_{k},P^{\prime}_{k}=\left(\frac{{P_{k}}}{\sum_{j=1}^{K}{P_{j}}}\right)$ (2)
where $P_{k}$ and $P^{\prime}_{k}$ mean the probability and normalized probability of token $k$.
@@ 3.2 Sentence-level Self-Reflection
Different prompts can significantly affect outputs of LLMs (Kung et al., 2023; Peng et al., 2023b), introducing uncertainty into IT data evaluation at the sentence level. To make better use of this uncertainty to bolster the reliability of our method, we implement sentence-level self-reflection (Sentence-R). Building upon Token-R, we devise $K$ semantically similar rating prompts$\{RP_{0},RP_{1},\ldots,RP_{K}\}$ to obtain a series of quality scores $\{S^{token}_{0},S^{token}_{1},\ldots,S^{token}_{K}\}$ based on a given sample $S$. We calculate the average of these scores to represent the overall quality of sample $S$, because of the importance of incorporating assessments from diverse prompts. Additionally, we use the standard deviation to quantify the LLMs’ uncertainty to rating prompt; a higher standard deviation suggests greater sensitivity to prompt variation, while a lower standard deviation indicates more consistent and confident quality ratings by LLMs (Zhou et al., 2020). By integrating a holistic sample evaluation with the quantification of model uncertainty, we derive the sentence-level score $S^{sent}$, offering a more nuanced and reliable measure of IT data quality.
$S^{sent}=\frac{\textbf{Avg}\{S^{token}_{i}\}^{K}_{i=1}}{1+\alpha\times\underbrace{\textbf{Std}\{S^{token}_{i}\}^{K}_{i=1}}_{Uncertainty}}$ (4)
@@ 3.3 Model-level Self-Reflection
A sample affirmed by multiple foundation models can truly be deemed as high-quality. Different foundation models have different quality assessments of the sample, which introduce model-level uncertainty. To maximize the utilization of this uncertainty, we introduce model-level self-reflection (Model-R). This strategy leverages the capabilities of existing open-source models without the need for additional resources or the complexities associated with fine-tuning. However, the challenge lies in the diverse capabilities of various LLMs and determining how to reasonably combine their sample evaluation based on their performance. It is widely acknowledged that the capabilities of LLMs tend to increase with their parameter count (Hendrycks et al., 2021). Thus, we suggest using the parameter count of LLMs as an initial metric for assessing their capabilities to properly weight sample quality scores. Given $N$ foundation models with parameter counts $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ and their respective sentence-level scores for a sample $S$ being $\{S^{sent}_{0},S^{sent}_{1},\ldots,S^{sent}_{N}\}$, we formulate the model-level score $S^{model}$ to reflect a comprehensive evaluation of sample quality.
@@
where $N$ means the number of the foundation models. By obtaining LLM parameters without resource expenditure, Model-R effectively allows us to employ more powerful foundation models, which is advantageous for selecting higher-quality data. Finally, we use $S^{model}$ as the final evaluation of sample $S$ in SelectIT. The higher $S^{model}$, the better sample quality. We sort the samples in descending order based on their $S^{model}$ and then select the top-ranked samples as high-quality data.
@@ 3.4 Selective Alpaca
We apply SelectIT to the widely-used Alpaca-GPT4 (Peng et al., 2023a). Specifically, we use the most popular LLaMA-2 (7B, 13B, 70B) as our foundation models and set the hyper-parameters $\alpha=0.2$ and $K=5$, which decides the range of LLMs rating in Token-R and the number of rating prompts in Sentence-R. We finally select the top 20%, a total of 10.4K pairs as the high-quality data and obtain a curated IT dataset called Selective Alpaca.
@@ 4.1 Setups
To gain a more comprehensive understanding of the capabilities of LLMs, we evaluate our approach in diverse downstream tasks (Wang et al., 2024; Ivison et al., 2023). Factual knowledge: We use the Massive Multitask Language Understanding dataset (MMLU (Hendrycks et al., 2021)) to assess the factual knowledge of LLMs and report 5-shot results. Reasoning: We evaluate the reasoning abilities of LLMs using two widely utilized datasets: the Grade School Math dataset (GSM (Cobbe et al., 2021)) and Big-Bench-Hard (BBH (Suzgun et al., 2022)) with the CoT setting (Wei et al., 2022b). Multilinguality: we assess this ability by TyDiQA, a multilingual question-answering benchmark that encompasses 11 diverse languages, with the gold-passage setup. Coding: We evaluate this ability using the HumanEval dataset (Chen et al., 2021) and report pass@10 results with a temperature of 0.8. Open-ended generation: We utilize AlpacaEval (Dubois et al., 2023), which employs GPT-4 to effectively assess model outputs. This can evaluate whether the text produced by LLMs aligns with humans.
@@
We compare with the following baselines:
@@
  • LIMA (Zhou et al., 2023) primarily consists of 1000 manually crafted high-quality instructional data, which can better stimulate the alignment capability of LLMs.
  • AlpaGasus (Chen et al., 2024) involves utilizing the robust ChatGPT to score and select data from the original Alpaca-GPT4 dataset.
@@
ID   System External MMLU BBH GSM TydiQA CodeX AE Overall
Model Data AVG $\Delta$ ($\uparrow$)
Base Model: LLaMA-2-7B Implemented Existing Method
1    Alpaca-GPT4 46.5 38.4 15.0 43.4 26.8 34.2 34.1 -
2    LIMA 45.4 37.5 14.3 45.1 24.6 33.1 33.3 -0.7
3    1 + AlpaGasus 45.9 39.0 14.5 46.4 27.5 35.4 34.8 +0.7
4    1 + Q2Q 46.9 39.4 15.3 46.7 28.2 35.7 35.4 +1.3
5    1 + Instruction Mining 47.0 39.6 16.5 47.1 28.6 34.4 35.5 +1.5
[…]
6    1 + Token-R 46.8 36.5 14.5 44.6 28.9 35.5 34.5 +0.4
7    1 + Sentence-R 46.9 38.1 16.1 48.4 26.9 35.3 35.3 +1.2
8    1 + Model-R 47.3 37.4 16.1 45.3 28.4 35.8 35.1 +1.0
[…]
9    SelectIT (6 + 7 + 8) 47.4 40.6 16.8 47.4 29.4 35.7 36.2 +2.2
Base Model: LLaMA-2-13B Implemented Existing Method
10    Alpaca-GPT4 55.7 46.6 30.5 48.1 40.8 46.5 44.7 -
11    LIMA 54.6 45.3 30.5 51.1 34.1 42.6 43.0 -1.7
12   10 + AlpaGasus 54.1 47.3 31.5 50.6 41.3 46.3 45.2 +0.5
13   10 + Q2Q 55.3 48.5 32.0 50.8 41.3 47.3 45.9 +1.2
14   10 + Instruction Mining 54.1 47.3 32.5 52.6 43.3 48.3 46.3 +1.6
[…]
15   10 + Token-R 55.3 47.3 30.5 51.3 39.8 46.2 45.1 +0.4
16   10 + Sentence-R 55.2 48.3 31.0 52.2 42.5 46.3 45.9 +1.2
17   10 + Model-R 55.1 47.5 31.5 52.3 40.2 46.1 45.5 +0.8
[…]
18   SelectIT (15 + 16 + 17) 55.7 48.9 33.0 54.1 42.2 48.8 47.1 +2.4
Table 1: Overall results on IT. “CodeX” and “AE” mean HumanEval and AlpacaEval benchmarks. All the scores are averages of three independent runs with different random seeds.